OpenStudy (anonymous):

I need to find the four arithmetic means between -21 and -36

OpenStudy (solomonzelman):

\(\large\color{darkblue}{ \rm -21,~~\text{_},~~\text{_},~~\text{_},~~\text{_},~~36. }\) \(\large\color{darkblue}{ \rm -21+5d=-36. }\) \(\large\color{darkblue}{ \rm 5d=-15. }\) \(\large\color{darkblue}{ \rm d=-3. }\) \(\large\color{darkblue}{ \rm 21,\color{red} { ~-24,~-27,~-30,~-33,} ~-36 }\)

OpenStudy (anonymous):

:O THANKYOUU! pleaseee help me with another onee!!

OpenStudy (anonymous):

I need to find the four arithmetic means between 100 and 135

OpenStudy (solomonzelman):

can you try it based on the previous problem ?

OpenStudy (anonymous):

100+5d=135?

OpenStudy (solomonzelman):

100, ` ___ ` ` ___ ` ` ___ ` ` ___ ` 135 yes... !

OpenStudy (solomonzelman):

Find the d

OpenStudy (solomonzelman):

(d is the difference btw the terms )

OpenStudy (solomonzelman):

`100+5d=135 ` 1) subtract 100 from both sides 2) divide both sides by 5

OpenStudy (anonymous):

5d=35?

OpenStudy (solomonzelman):

yes and divide both sides by 5 ?

OpenStudy (solomonzelman):

d = ?

OpenStudy (anonymous):

d=7?

OpenStudy (solomonzelman):

So can you fill in the blanks ? 100, `___` `___` `___` ___` 135

OpenStudy (anonymous):

107,114,121,128?

OpenStudy (solomonzelman):

Yeeeah !!

OpenStudy (anonymous):

THANKYOUUUU so muchh!

OpenStudy (solomonzelman):

So when you have N arithmetic means between A and B you say A+Nd=C and solve for d, knowing the number of means N and knowing the number A and B makes sense ?

OpenStudy (anonymous):

it does thankyou soo much! would you help me with one more? please:)?

OpenStudy (solomonzelman):

Sure :)

OpenStudy (anonymous):

What is the 20th term of the following arithmetic sequence? 1/2, 1, 3/2, 2

OpenStudy (solomonzelman):

1 - 1/2 = ? (this would be the difference - what you add to each term, to obtain the next term. ) (for the beginning )

OpenStudy (solomonzelman):

So ` d = ½ `

OpenStudy (anonymous):

ok :)?

OpenStudy (solomonzelman):

now use the formula, \(\large\color{darkblue}{ \rm a_1+d(n-1)=a_n }\) In your case, \(\large\color{darkblue}{ \rm \frac{1}{2}+ \frac{1}{2}(20-1)=a_{20} }\) Makes sense ?

OpenStudy (anonymous):

yes totally! so..

OpenStudy (solomonzelman):

Can you find the \(\large\color{black}{ \rm a_{20} }\) ?

OpenStudy (solomonzelman):

½ + ½(20-1) = A 20 ½ + ½(19) = A 20 ½(1) + ½(19) = A 20 ½(20) = A 20 10=A 20

OpenStudy (anonymous):

1(19)=a20?

OpenStudy (solomonzelman):

See how I am getting A20 ?

OpenStudy (anonymous):

I see now, ok I think I got it!

OpenStudy (anonymous):

THANKYOUUUU!!!! :)

OpenStudy (solomonzelman):

Anytime :)